OpenStudy (anonymous):

A system of equations is shown below: -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

Part A: equation (1) -3x + 7y = -16 equation (2)-9x + 5y = 16 replace (2) by adding (2) to two times (1): 2 times (1): -6x + 14y = -32 Add (2): -9x + 5y = 16 -------------------- -15x + 19y = -16 So the second pair of equations are: -3x + 7y = -16 (1) -15x + 19y = -16 (3)

OpenStudy (anonymous):

I just need help with Part B

OpenStudy (anonymous):

@undeadknight26 @itsbribro

OpenStudy (anonymous):

i think it would be solve for 1 and 2 x,y And then solve for 1 and 3 x,y

OpenStudy (anonymous):

@KlOwNlOvE i need you help

OpenStudy (anonymous):

@Catlover5925 @dtuniyants @dan815

OpenStudy (anonymous):

@perl

OpenStudy (camerondoherty):

I'm not a dude -.-

OpenStudy (anonymous):

okay sorry will you please help me tho.

OpenStudy (camerondoherty):

And I have no idea how to do this, sorry

OpenStudy (anonymous):

no hold on

OpenStudy (anonymous):

i just need help finding x and y for -3x+y=-16 -9x+5y=16 -15x+19x=-16

OpenStudy (anonymous):

-3x+7y=-16

OpenStudy (anonymous):

@perl

OpenStudy (perl):

it would be easier if you redid part a)

OpenStudy (anonymous):

why was it wrong?

OpenStudy (anonymous):

but if you will help me through it i will.

OpenStudy (perl):

its not wrong, but you didn't solve for x and y

OpenStudy (anonymous):

okay so we will restart where do you want to start?

OpenStudy (perl):

actually lets keep your original work

OpenStudy (perl):

brb

OpenStudy (anonymous):

okay

OpenStudy (perl):

so we want to show that the 'new system of equations' has the same solution as the 'old system of equations'

OpenStudy (anonymous):

yes

OpenStudy (perl):

to show that, lets plug in the solution into both system

OpenStudy (anonymous):

first can you make sure my part A is correct?

OpenStudy (perl):

ok

OpenStudy (perl):

i had to reboot my computer , this website is super glitchy

OpenStudy (anonymous):

yeah okay so can you check then go to part B?

OpenStudy (perl):

part b) show that system : -3x + 7y = -16 -9x + 5y = 16 has same solution as system : -3x + 7y = -16 -15x + 19y = -16

OpenStudy (anonymous):

yes

OpenStudy (perl):

ok plug in x = -4, y = -4 (that is the solution for the first system )

OpenStudy (anonymous):

so where do i plug it in?

OpenStudy (perl):

for x and y

OpenStudy (anonymous):

for all 4 of them?

OpenStudy (anonymous):

12+(-28) 36+(-20) 60+(-76)

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

is that right? i checked and it was

OpenStudy (perl):

part b) show that system : -3x + 7y = -16 -9x + 5y = 16 has same solution as system : -3x + 7y = -16 -15x + 19y = -16 Plug in (x,y) = (-4,-4) system 1: -3(-4) + 7(-4) = 12 - 28= -16 -9(-4) + 5(-4) = 36 - 20 = 16 system 2: -3(-4) + 7(-4) = 12-28 = -16 -15(-4) + 19(-4) = 60 -76= -16 so (-4,-4) is the solution for both systems. therefore the systems of equations are equivalent

OpenStudy (anonymous):

Thanks so much